Question 122 points total). Your sister is starting to plan for your nieces college expenses and has asked for your advice.


Do not use excell or any software. Use formulas/equations, and show calculations. Thank you.
Answer #1

Deposit = $ 5,000, Growth, G = $ 1,500

Number of deposits = 12

Interest rate = 6%

F = 5,000(F/A, 6%, 12) +1,500(F/G,6%, 12)

= 1.0612_1 F= 5,000 x = 1.0612 – 0.06 x 12 - 1 0,062 0.06 +1, 500x

S F =5,000 x 16.8699 +1,500 x 81.16568

=> F= 84,349.71 + 121,748.53

Future value = $ 206,098.24

B. The fees will increase at a rate of 2.5% every year. We can determine the present value of the fees. Assume the first fees paid is $ A.

PV = A(P/G, 2.5%,6%,5)

A — PV = 2,06 -0.025 * [1 – (1 02 PV = a 0.06 – 0.025 1.025 1.06

PV = AX 4.415599

This must be equal to $ 206,098.24

A x 4.415599 = 206.098.24

= A= 206,098.24 4.415599

A ~$ 46,675

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me, please help. Thank you.